Determine the equation of a line passing through (0,4) with a slope of 1/3. Express this equation using the standard form.
Answer by mananth(16946)

(0,4) with a slope of 1/3
standard equation = y=mx+c
4=1/3 *0+c
c=4
y=1/3 *x+4
y=x/3 +4